# Automatic Curriculum Expert Iteration for Reliable LLM Reasoning

### v1

@@ Abstract
Hallucinations (i.e., generating plausible but inaccurate content) and laziness (i.e. excessive refusals or defaulting to “I don’t know”) persist as major challenges in LLM reasoning. Current efforts to reduce hallucinations primarily focus on factual errors in knowledge-grounded tasks, often neglecting hallucinations related to faulty reasoning. Meanwhile, some approaches render LLMs overly conservative, limiting their problem-solving capabilities. To mitigate hallucination and laziness in reasoning tasks, we propose Automatic Curriculum Expert Iteration (Auto-CEI) to enhance LLM reasoning and align responses to the model’s capabilities–assertively answering within its limits and declining when tasks exceed them. In our method, Expert Iteration explores the reasoning trajectories near the LLM policy, guiding incorrect paths back on track to reduce compounding errors and improve robustness; it also promotes appropriate “I don’t know” responses after sufficient reasoning attempts. The curriculum automatically adjusts rewards, incentivizing extended reasoning before acknowledging incapability, thereby pushing the limits of LLM reasoning and aligning its behaviour with these limits. We compare Auto-CEI with various SOTA baselines across logical reasoning, mathematics, and planning tasks, where Auto-CEI achieves superior alignment by effectively balancing assertiveness and conservativeness.
@@ 1 Introduction
Hallucination is a long-standing issue in large language model (LLM) research, which refers to the phenomenon where LLM-generated content appears plausible but is actually nonsensical or inaccurate, often misleading humans by seeming deceptively credible (Ji et al., 2023). It is more evident when solving complex reasoning problems beyond their capability, in which LLM tends to fake evidence or logic to answer the questions assertively. The reason for LLM’s hallucinations, overall, is a misalignment of its real capability and its behaviours: LLM should not behave overly assertively or confidently when facing unfamiliar (Kang et al., 2024a) and difficult problems (Dziri et al., 2024). Instead, a preferred behaviour of LLMs is to acknowledge their limitations by responding with “I don’t know.” However, we must also avoid the LLM from defaulting to degenerate responses like “I don’t know,” especially when the question is within its capability. A reliable LLM should strike a balance between maximizing performance and avoiding hallucination.
@@ 2 Related work
Reinforcement learning for LLM reasoning. Reinforcement learning has been widely used in the alignment stage of LLM training to enhance their capabilities. The research in this area can be broadly divided into two aspects: (1) RL algorithms; (2) reward construction. Conventional advanced deep RL algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), have been extensively applied in the training process of modern LLMs, helping to improve both format and reasoning abilities (Achiam et al., 2023; Team et al., 2023; Dubey et al., 2024). However, due to the introduction of the value network, the computational overhead of these algorithms is extremely high, often making them unaffordable for academic settings where resources are limited. More recently, expert iteration (also known as rejection sampling or reward-ranked fine-tuning) (Dong et al., 2023; Gulcehre et al., 2023), preference-based RL (Rafailov et al., 2024; Xiong et al., 2024a), and REINFORCE-type RL (Williams, 1992; Shao et al., 2024; Li et al., 2023; Ahmadian et al., 2024) have emerged as more efficient alternatives for the alignment stage, especially when operating within a fixed budget. These methods have demonstrated significant improvements over supervised fine-tuning in reasoning tasks (Aksitov et al., 2023; Gou et al., 2023; Havrilla et al., 2024; Shao et al., 2024; Dong et al., 2024; Trung et al., 2024; Xiong et al., 2024b). Additionally, Lightman et al. (2023) introduced process supervision to refine the reward signal for each reasoning step. Compared to traditional outcome-based supervision, process supervision has been shown to yield substantial gains (Lightman et al., 2023; Wang et al., 2024; Shao et al., 2024). Despite the impressive progress in this area, most existing works focus on maximizing the upper bound of reasoning ability. In contrast, our algorithm aims to develop a more balanced model incorporating self-awareness into the reasoning procedure.
@@ 4.2 Expert Iteration
The reward function helps us build the expert, enabling us to improve LLM reasoning performance over time. We resample the responses according to the rewards to build the expert. We first use LLM to generate $K$ samples $\{y_{i}\}_{i=1}^{K}$ for each of the questions in the training dataset $D$. Then, we use the reward function to get a reward value of each sample $r_{1},r_{2},\ldots,r_{K}$. Over that reward, we employ the temperature-scaled softmax function to construct a new distribution: $p^{\prime}(y_{i})=\exp(r_{i}/\tau)/\sum_{i}\exp(r_{i}/\tau)$ and subsequently use $p^{\prime}(y_{i})$ to conduct resampling and generate $N$ new samples $\{y^{\prime}_{i}\}_{i=1}^{K}$. $\tau$ is the temperature of resampling, the same as the overall accuracy of the initial SFT model finetuned by $D$. If the SFT model has high overall accuracy, we will encourage LLM to explore more randomly; otherwise, the responses with higher rewards will be sampled more densely. $\tau$ will be capped in a range $[0.4,0.7]$ to avoid extreme cases. Thus, the distribution is reshaped to encourage more instances that produce correct answers and produce more reasoning steps before saying “I don’t know.” Since resampling also kills instances that waste some of the exploration data, we concatenate the $\{y_{i}\}_{i=1}^{K}$ and $\{y^{\prime}_{i}\}_{i=1}^{K}$ together to form the new training dataset. Lastly, we removed the answers that had $-1$ rewards.
@@ 4.3 Curriculum update
3:   Update $c_{1}$ in reward function $R$triangleright Equ 1
@@ 5.1 Experimental Setup
Benchmarks. To demonstrate the effectiveness of Auto-CEI in reasoning tasks, we select BoardgameQA (Kazemi et al., 2024), MATH (Hendrycks et al., 2021), and Blocksworld (Valmeekam et al., 2023) as benchmarks, spanning from logical and mathematical reasoning to planning. They have various domains and complexities. We briefly introduce the benchmarks and report our detailed experimental settings in the Appendix B.
  • BoardgameQA is one of the latest benchmarks for logical reasoning. The data is synthesised from formal logical reasoning rules and clauses. The problem provides a self-contained context, pieces of evidence, reasoning rules, and questions. Given the evidence and reasoning rules, LLM must decide if the queried clause is true, false, or unknown. Moreover, BoardgameQA has cases where some evidence might contradict others. To deal with this case, BoardgameQA also provides preferences among the reasoning rules, meaning that the conclusions drawn from the preferred rules have higher priority. In addition, since BoardgameQA does not provide CoT data for unknown questions, we use GPT-4 to generate CoT reasoning trajectories for unknown problems for the training dataset. Thus, the training data is consistent. The prompt for data generation is in the Appendix B.1.
@@
Implementation details. We use Llama-3.1-8B-instruct (Dubey et al., 2024) as the backbone model and use Lora ($r=128,\alpha=64$) to fine-tune. The experiments are conducted in a server with 8$\times$×Nvidia A100 (40GB) GPUs. We use DeepSpeed Stage 2 to conduct the training. See Section C in Appendix for more details about the Auto-CEI setup.
@@ 6 Conclusion
This work focuses on challenging reasoning problems, where complexity can be estimated by the number of reasoning steps. For simpler problems with shorter reasoning paths or well-studied datasets (e.g., GSM8K), R-Tuning reliably performs well, leaving limited scope for further optimization. This is partly due to the reduced impact of compounding errors in these easier tasks. We also found that the model might decline to answer even if its response is correct. Thus, further adjustment and fine-tuning should be applied to solve this issue, and our method should be able to be extended accordingly as well. In addition, our method might require a long time for optimisation as it requires running Expert Iteration multiple times. Nonetheless, compared to the cost of pre-training, the overhead of the iterative post-training algorithm remains worthwhile for achieving better alignments.
Reproducibility Statement
@@
In Section 4, we provide pseudocode for our algorithm and explain the details of each part. We provide sufficient details in Appendix Section C for our implementation of Auto-CEI and training settings. For our experiment benchmarks, we introduce our benchmark selection and metric computation in Section 5.1 and provide the link to download the dataset or the source code from the published papers to generate the dataset, the format of the dataset, and the dataset size we choose in Appendix Section B. We also provide the baseline methods’ details, implementation and training in Appendix Section D.
Codes are available upon request and can be shared privately before the public release.
Appendix A Improvement in Expert Iteration
@@
Appendix B Experimental Settings
B.1 BoardgameQA
For boardgameQA, we use the main dataset to do fine-tuning and evaluation  [^1] The dataset is downloaded from https://storage.googleapis.com/gresearch/BoardgameQA/
BoardgameQA.zip
. We take the dataset under the folder BoardgameQA-Main-depth1, BoardgameQA-Main-depth2 and BoardgameQA-Main-depth3.
We show one example of the problem in BoardgameQA in Listing 1. For more details about the dataset itself, please refer to the paper by Kazemi et al. (2024).
@@ B.1 BoardgameQA
B.2 MATH
We download the MATH dataset via Hugging Face at https://huggingface.co/datasets/
hendrycks/competition_math
. It automatically divide the dataset into training, validation, and test set. We used all the training data in our experiment and evaluated it using all validation and testing sets to check if the final answer was the same as the labelled data. For more details about the dataset itself, please refer to the paper by Hendrycks et al. (2021). One example format of the dataset is shown in the Listing 3.
@@ B.2 MATH
B.3 Blocksworld
The Blocksworld dataset can be generated using the code in the GitHub repository by Valmeekam et al. (2023): https://github.com/karthikv792/LLMs-Planning. For our case, we generate domains from 4 blocks to 6 blocks, and randomly samples 500 data points for training. We randomly sample 500 data points for validation and testing sets whose optimal solution length (i.e., ground truth plan) is no longer than ten steps. We uniformly sample the tasks according to the ground truth lengths to form the testing set (i.e., 100 two-step tasks, 100 four-step tasks, …, and 100 ten-step tasks).
@@ B.3 Blocksworld
Appendix C Implementation details
C.1 Auto-CEI Setting
@@
We use Llama-3.1-8B-instruct (Dubey et al., 2024) as the backbone model and use Lora ($r=128,\alpha=64$) to fine-tune. The Llama-3.1-8B-instruct model is downloaded at https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct. We use the Fastchat template designed for Llama-3.1-8B-Instruct. The template is demonstrated in the Listing 5. We fill in the Prompt by our query and train the LLM to respond by Completion.
@@
C.2 Training details
The experiments are conducted in a server with 8$\times$×Nvidia A100 GPUs. We use DeepSpeed Stage 2 to conduct the training  [^2] The configuration of DeepSpeed Stage 2 can be found at https://github.com/huggingface/transformers/blob/
main/tests/deepspeed/ds_config_zero2.json
. During the SFT stage, we choose the batch size = 256, and epoch number is 5 for initial SFT process, and 2 for SFT in R-Tuning and EI process. The learning rate of SFT is $1\times 10^{-4}$. We use the HuggingFace SFT trainer [^3] https://huggingface.co/docs/trl/v0.11.1/en/sft_trainer#trl.SFTTrainer to conduct the SFT training automatically.
Appendix D Details of baseline methods
D.1 R-Tuning
The setting is the same as Auto-CEI in the initialisation part of Section C.1.
D.2 RLKF

### v2

@@ Abstract
Hallucinations (i.e., generating plausible but inaccurate content) and laziness (i.e. excessive refusals or defaulting to “I don’t know”) persist as major challenges in LLM reasoning. Current efforts to reduce hallucinations primarily focus on factual errors in knowledge-grounded tasks, often neglecting hallucinations related to faulty reasoning. Meanwhile, some approaches render LLMs overly conservative, limiting their problem-solving capabilities. To mitigate hallucination and laziness in reasoning tasks, we propose Automatic Curriculum Expert Iteration (Auto-CEI) to enhance LLM reasoning and align responses to the model’s capabilities–assertively answering within its limits and declining when tasks exceed them. In our method, Expert Iteration explores the reasoning trajectories near the LLM policy, guiding incorrect paths back on track to reduce compounding errors and improve robustness; it also promotes appropriate “I don’t know” responses after sufficient reasoning attempts. The curriculum automatically adjusts rewards, incentivizing extended reasoning before acknowledging incapability, thereby pushing the limits of LLM reasoning and aligning its behaviour with these limits. We compare Auto-CEI with various SOTA baselines across logical reasoning, mathematics, and planning tasks, where Auto-CEI achieves superior alignment by effectively balancing assertiveness and conservativeness.  [^1] The code is available at https://github.com/SalesforceAIResearch/Auto-CEI.
@@ 1 Introduction
Hallucination is a long-standing issue in large language model (LLM) research, which refers to the phenomenon where LLM-generated content appears plausible but is actually nonsensical or inaccurate, often misleading humans by seeming deceptively credible (Ji et al., 2023). It is more evident when solving complex reasoning problems beyond their capability, in which LLM tends to fake evidence or logic to answer the questions assertively. The reason for LLM’s hallucinations, overall, is a misalignment of its real capability and its behaviours: LLM should not behave overly assertively or confidently when facing unfamiliar (Kang et al., 2024a) and difficult problems (Dziri et al., 2024). Instead, a preferred behaviour of LLMs is to acknowledge their limitations by responding with “I don’t know.” However, we must also prevent the LLM from defaulting to degenerate responses like ”I don’t know,” especially when the question is within its capability. A reliable LLM should strike a balance between maximizing performance and avoiding hallucination.
@@ 2 Related work
Reinforcement learning for LLM reasoning. Reinforcement learning has been widely used in the alignment stage of LLM training to enhance their capabilities. The research in this area can be broadly divided into two aspects: (1) RL algorithms; (2) reward construction. Conventional advanced deep RL algorithms, such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), have been extensively applied in the training process of modern LLMs, helping to improve both format and reasoning abilities (Achiam et al., 2023; Team et al., 2023; Dubey et al., 2024). However, due to the introduction of the value network, the computational overhead of these algorithms is extremely high, often making them unaffordable for academic settings where resources are limited. More recently, expert iteration (also known as rejection sampling or reward-ranked fine-tuning) (Dong et al., 2023; Gulcehre et al., 2023; Zelikman et al., 2022; Hoffman et al., 2024), preference-based RL (Rafailov et al., 2024; Xiong et al., 2024a), and REINFORCE-type RL (Williams, 1992; Shao et al., 2024; Li et al., 2023; Ahmadian et al., 2024) have emerged as more efficient alternatives for the alignment stage, especially when operating within a fixed budget. These methods have demonstrated significant improvements over supervised fine-tuning in reasoning tasks (Aksitov et al., 2023; Gou et al., 2023; Havrilla et al., 2024; Shao et al., 2024; Dong et al., 2024; Trung et al., 2024; Xiong et al., 2024b). Additionally, Lightman et al. (2023) introduced process supervision to refine the reward signal for each reasoning step. Compared to traditional outcome-based supervision, process supervision has been shown to yield substantial gains (Lightman et al., 2023; Wang et al., 2024a; Shao et al., 2024). Despite the impressive progress in this area, most existing works focus on maximizing the upper bound of reasoning ability. In contrast, our algorithm aims to develop a more balanced model incorporating self-awareness into the reasoning procedure.
@@ 4.2 Expert Iteration
The reward function helps us build the expert, enabling us to improve LLM reasoning performance over time. We resample the responses according to the rewards to build the expert. We first use LLM to generate $K$ samples $\{y_{i}\}_{i=1}^{K}$ for each of the questions in the training dataset $D$. Then, we use the reward function to get a reward value of each sample $r_{1},r_{2},\ldots,r_{K}$. Over that reward, we employ the temperature-scaled softmax function to construct a new distribution: $p^{\prime}(y_{i})=\exp(r_{i}/\tau)/\sum_{j}\exp(r_{j}/\tau)$ and subsequently use $p^{\prime}(y_{i})$ to conduct resampling and generate $N$ new samples $\{y^{\prime}_{i}\}_{i=1}^{K}$. $\tau$ is the temperature of resampling, the same as the overall accuracy of the initial SFT model finetuned by $D$. If the SFT model has high overall accuracy, we will encourage LLM to explore more randomly; otherwise, the responses with higher rewards will be sampled more densely. $\tau$ will be capped in a range $[0.4,0.7]$ to avoid extreme cases. Thus, the distribution is reshaped to encourage more instances that produce correct answers and produce more reasoning steps before saying “I don’t know.” Since resampling also kills instances that waste some of the exploration data, we concatenate the $\{y_{i}\}_{i=1}^{K}$ and $\{y^{\prime}_{i}\}_{i=1}^{K}$ together to form the new training dataset. Lastly, we removed the answers that had $-1$ rewards.
@@ 4.3 Curriculum update
3:   Update $c_{1}$ in reward function $R$triangleright Eq 1
@@ 5.1 Experimental Setup
Benchmarks. To demonstrate the effectiveness of Auto-CEI in reasoning tasks, we select BoardgameQA (Kazemi et al., 2024), MATH (Hendrycks et al., 2021), and Blocksworld (Valmeekam et al., 2023) as benchmarks, spanning from logical and mathematical reasoning to planning. They have various domains and complexities. We briefly introduce the benchmarks and report our detailed experimental settings in the Appendix F.
  • BoardgameQA is one of the latest benchmarks for logical reasoning. The data is synthesised from formal logical reasoning rules and clauses. The problem provides a self-contained context, pieces of evidence, reasoning rules, and questions. Given the evidence and reasoning rules, LLM must decide if the queried clause is true, false, or unknown. Moreover, BoardgameQA has cases where some evidence might contradict others. To deal with this case, BoardgameQA also provides preferences among the reasoning rules, meaning that the conclusions drawn from the preferred rules have higher priority. In addition, since BoardgameQA does not provide CoT data for unknown questions, we use GPT-4 to generate CoT reasoning trajectories for unknown problems for the training dataset. Thus, the training data is consistent. The prompt for data generation is in the Appendix F.1.
@@
Implementation details. We use Llama-3.1-8B-instruct (Dubey et al., 2024) as the backbone model and use Lora ($r=128,\alpha=64$) to fine-tune. The experiments are conducted in a server with 8$\times$×Nvidia A100 (40GB) GPUs. We use DeepSpeed Stage 2 to conduct the training. See Section G in Appendix for more details about the Auto-CEI setup.
@@ 6 Conclusion
This work focuses on challenging reasoning problems, where complexity can be estimated by the number of reasoning steps. For simpler problems with shorter reasoning paths or well-studied datasets (e.g., GSM8K), R-Tuning reliably performs well, leaving limited scope for further optimization. This is partly due to the reduced impact of compounding errors in these easier tasks. We also found that the model might decline to answer even if its response is correct. Thus, further adjustment and fine-tuning should be applied to solve this issue, and our method should be able to be extended accordingly as well. In addition, our method might require a long time for optimisation as it requires running Expert Iteration multiple times. Nonetheless, compared to the cost of pre-training, the overhead of the iterative post-training algorithm remains worthwhile for achieving better alignments. Our primary goal was to establish a proof-of-concept using straightforward, reproducible metrics common in curriculum learning. While we focused on the number of reasoning steps as the core metric, the framework can accommodate more sophisticated approaches. Future work could incorporate advanced metrics such as LLM-predicted statistics (Wang et al., 2024b; Damani et al., 2024) for fine-grained curriculum learning. The foundational nature of Auto-CEI makes it amenable to such extensions while maintaining its core benefits of automation and scalability.
Reproducibility Statement.
@@
In Section 4, we provide pseudocode for our algorithm and explain the details of each part. We provide sufficient details in Appendix Section G for our implementation of Auto-CEI and training settings. For our experiment benchmarks, we introduce our benchmark selection and metric computation in Section 5.1 and provide the link to download the dataset or the source code from the published papers to generate the dataset, the format of the dataset, and the dataset size we choose in Appendix Section F. We also provide the baseline methods’ details, implementation and training in Appendix Section H. Codes are available.
Acknowledgment
We thank all Program chairs, Area chairs, and Reviewers for helpful discussion.
Appendix A Difficulty and response length
The underlying assumption of our paper is that difficult problems require longer responses to get the answer. We use the MATH dataset, which has manually labelled difficulties, to verify the hypothesis. We measure the steps of reasoning with respect to the labelled difficulty, shown in Fig 4. The result suggests that the reasoning length and manually labelled difficulty are correlated with statistical significance. The scale looks different from Fig 3 in the paper, as the reasoning length distribution is a long tail distribution. The distribution is visualised in Fig 5.
Human-labelled difficulty often diverges from computational measures, with LLM-perceived difficulty better aligning with reasoning steps (Pearson’s $r$ = 0.349, $p<$ 1e-142). Despite the noise in the MATH dataset’s non-standardized steps, Auto-CEI performs robustly, effectively balancing assertiveness and conservativeness.
Figure 4: Correlation between reasoning length and labeled difficulty (MATH dataset). For each difficulty level, the mean reasoning length is plotted with error bars representing the standard error of the estimator.
Figure 5: Histogram of reasoning lengths generated by Auto-CEI for the MATH dataset.
Some difficult problems might have short optimal solutions. However, this doesn’t mean the computational complexity of that problem is low. Take NP-complete problems as an example. Searching for the solution may require exponential time (in the worst case), but verifying the solution only takes polynomial time, as the optimal solution is much shorter than the search process. While the question of whether verification is inherently simpler than finding a solution (NP ?= P) remains open, our hypothesis relies on human intuitive common sense: more difficult problems generally require higher computational complexity to arrive at correct solutions, which is also the case for real-world problems and solutions.
Appendix B Hyperparameter tuning
The mitigation of hallucination and laziness contradict each other. Optimizing precision makes the LLM respond “I don’t know” most of the time; minimising the refusal rate makes the LLM try to make an assertive response all the time. Because of this, our objective is not to maximise accuracy but to achieve the sweet spot of balance between assertiveness and conservativeness.
The hyper-parameter does not affect the training process but determines when to stop the Auto-CEI. Empirically, the user can start an initial hyper-parameter. After it terminates, if the user is unsatisfied with the result, they can further update the hyper-parameter and continue the training process. The previous models can be reused so that no extra computation is involved. One example is shown in Fig 6.
Figure 6: The value of objective function $f$ changes according to the curriculum and hyper-parameter $\lambda$. The red dot denotes the highest value of $f$ in the curve.
Appendix C Refusal accuracy
We evaluate if the language model makes a refusal response even if its answer is correct. Thus, we define the refusal accuracy (IDK Acc) below.
$P_{\mathrm{IDKAcc}}=\frac{\text{Number of refusal response with correct answer}}{\text{Number of refusal response}}$ (3)
The refusal accuracy measures the proportion of the correct answer in the refusal responses. A lower accuracy indicates a better self-awareness in correctness, while a higher correctness means the language model is over-conservative. The refusal accuracy, compared with the precision, is shown in Table 3.
Table 3: Refusal accuracy v.s. Precision. Pre: precision (%); IDK Acc: refusal accuracy (%).
Boardgame QA MATH BlocksWorld
Pre IDK Acc Pre IDK Acc Pre IDK Acc
SFT + R-Tuning 80.36 19.62 60.67 14.27 90.69 32.44
EI + R-Tuning 80.77 16.67 55.80 12.74 93.95 41.21
RLKF 54.17 21.48 42.19 13.49 38.08 24.47
Auto-CEI 84.52 12.25 55.63 9.89 91.53 27.02
The result suggests that Auto-CEI’s refusal responses have significantly lower accuracy. We also highlight that a model’s true capacity is inherently a hidden variable and can only be estimated indirectly, as there is no ground-truth label of which questions are truly beyond LLM’s capacity. 1 - (refusal accuracy) is a Monte-Carlo estimation of the model’s awareness in providing an inaccurate answer.
Appendix D Generalizability cross language models
We conduct additional experiments to test the generalizability of our method in other language models. We did experiments on Mistral-7B-Instruct-v3 on the BoardgameQA benchmark. The result is shown in Table 4. It provides further evidence of our method’s generalisability.
Table 4: Additional Experiment on Mistral-7B-Instruct-v3. Acc: accuracy (%); Pre: precision (%); IDK: refusal rate (%); IDK Acc: refusal accuracy (%).
Acc Pre IDK IDK Acc $f(\lambda=0.2)$
SFT 71.28
EI 75.54
SFT + R-Tuning 59.02 76.39 22.74 15.42 0.7656
EI + R-Tuning 61.62 80.39 23.35 14.23 0.7964
Auto-CEI 75.01 93.47 19.75 11.50 0.9083
Appendix E Improvement in Expert Iteration
@@
Appendix F Experimental Settings
F.1 BoardgameQA
For boardgameQA, we use the main dataset to do fine-tuning and evaluation  [^2] The dataset is downloaded from https://storage.googleapis.com/gresearch/BoardgameQA/BoardgameQA.zip. We take the dataset under the folder BoardgameQA-Main-depth1, BoardgameQA-Main-depth2 and BoardgameQA-Main-depth3. We show one example of the problem in BoardgameQA in Listing 1. For more details about the dataset itself, please refer to the paper by Kazemi et al. (2024).
@@ F.1 BoardgameQA
F.2 MATH
We download the MATH dataset via Hugging Face  [^3] Link: https://huggingface.co/datasets/hendrycks/competition_math. It automatically divide the dataset into training, validation, and test set. We used all the training data in our experiment and evaluated it using all validation and testing sets to check if the final answer was the same as the labelled data. For more details about the dataset itself, please refer to the paper by Hendrycks et al. (2021). One example format of the dataset is shown in the Listing 3.
@@ F.2 MATH
F.3 Blocksworld
The Blocksworld dataset can be generated using the code in the GitHub repository by Valmeekam et al. (2023) [^4] Link: https://github.com/karthikv792/LLMs-Planning.. For our case, we generate domains from 4 blocks to 6 blocks, and randomly samples 500 data points for training. We randomly sample 500 data points for validation and testing sets whose optimal solution length (i.e., ground truth plan) is no longer than ten steps. We uniformly sample the tasks according to the ground truth lengths to form the testing set (i.e., 100 two-step tasks, 100 four-step tasks, …, and 100 ten-step tasks).
@@ F.3 Blocksworld
Appendix G Implementation details
G.1 Auto-CEI Setting
@@
We use Llama-3.1-8B-instruct (Dubey et al., 2024) as the backbone model and use Lora ($r=128,\alpha=64$) to fine-tune  [^5] The Llama-3.1-8B-instruct model is downloaded at https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct.. We use the Fastchat template designed for Llama-3.1-8B-Instruct. The template is demonstrated in the Listing 5. We fill in the Prompt by our query and train the LLM to respond by Completion.
@@
G.2 Training details
The experiments are conducted in a server with 8$\times$×Nvidia A100 GPUs. We use DeepSpeed Stage 2 to conduct the training  [^6] The configuration of DeepSpeed Stage 2 can be found at https://github.com/huggingface/transformers/blob/main/tests/deepspeed/ds_config_zero2.json. During the SFT stage, we choose the batch size = 256, and epoch number is 5 for initial SFT process, and 2 for SFT in R-Tuning and EI process. The learning rate of SFT is $1\times 10^{-4}$. We use the HuggingFace SFT trainer [^7] https://huggingface.co/docs/trl/v0.11.1/en/sft_trainer to conduct the SFT training automatically.
Appendix H Details of baseline methods
H.1 R-Tuning
The setting is the same as Auto-CEI in the initialisation part of Section G.1.
H.2 RLKF